# 1Manifold Research 2Metarch.ai 3Georgia Tech

### v1

@@ IV-A1 Comparing General Performance
In evaluating the generalization capability of GPT 4o, GPT 4.1, OpenVLA, Pi0 Base, and Pi0 FAST, we observed universally poor performance on zero-shot out-of-distribution (OOD) tasks. The analysis using various performance metrics underscored significant limitations across all models.
As seen in Figure 1, The Brier Mean Absolute Error (MAE) scores revealed poor probability calibration, as scores approached the maximum value of 2 across all models and datasets. Specifically, GPT 4o, GPT 4.1, and Pi0 FAST consistently scored above 1.70, while OpenVLA generally exceeded 1.50. The subdatasets on which models exhibited the weakest performance were Starpilot for GPT 4o, Leaper for GPT 4.1, Miner for Pi0 FAST, and Ninja for OpenVLA.
Figure 1: Brier Mean Absolute Error scores across 4 models - GPT 4o, OpenVLA, GPT 4.1, and Pi0 FAST. Pi0 Base is a diffusion based model and can not be evaluated using Brier MAE due to lack of logits in its inference architecture. All models display Brier MAE close to 2 indicating poor performance.
Analyzing micro-level metrics, precision values were low across all models and subdatasets. OpenVLA achieved the highest micro precision at approximately 25 percent on the Coinrun dataset, whereas GPT 4o showed the lowest precision at 0.9 percent on Starpilot. Pi0 FAST reached a minimal precision of around 1 percent on Heist. Overall, as can be inferred from Figure 2, the performance hierarchy based on micro precision was: OpenVLA $>$ Pi0 Base $>$ GPT 4.1 $>$ GPT 4o $>$ Pi0 FAST.
@@
When considering micro precision without invalid predictions, as seen in Figure 3 significant improvements were noted for GPT 4o and Pi0 FAST, highlighting their potential if constrained or further fine-tuned to produce valid action outputs. GPT 4o’s precision notably improved, surpassing Pi0 Base in a few subdatasets such as Coinrun, Bossfight, and Dodgeball. Nevertheless, the adjusted performance order still positioned OpenVLA at the top, followed by Pi0 Base, Pi0 FAST, GPT 4o, and GPT 4.1.
Figure 3: Micro precision without considering invalids across all 5 models. GPT 4o and Pi0 FAST show a significant surge in precision once invalids are not considered.
@@
Macro precision analysis revealed that GPT 4o, OpenVLA, and Pi0 Base maintained similar performance levels across most datasets. As seen in Figure 4, conversely, GPT 4.1 and Pi0 FAST demonstrated significant variability, implying distinct preferences for specific task types. Despite these observations, macro precision remained generally low across all models, with a maximum slightly above 15 percent. Importantly, comparing its macro and micro precision performances, it is clear that OpenVLA’s performance heavily favored majority classes, while GPT 4o exhibited relatively stronger performance on minority classes.
@@
Figure 4: Macro precision across all 5 models. GPT 4.1 and Pi0 FAST show variability and preference towards specific subdatasets. OpenVLA shows reliability on majority classes whereas GPT 4o displays relatively stronger performance on minority classes.
@@
Macro recall, as seen in Figure 5 highlighted OpenVLA’s aggressive prediction strategy, resulting in consistently high recall but lower precision due to numerous false positives. GPT 4o, despite higher macro precision, showed substantially lower recall indicating pronounced class biases towards a few classes, and a potentially overly conservative nature about predicting certain classes compared to other classes. GPT 4.1 exhibited a relatively balanced approach, achieving both substantial recall and precision, thus indicating less bias towards specific classes. Pi0 Base displayed a moderately balanced performance without leading in any subdataset, whereas Pi0 FAST consistently showed low recall, reflecting challenges in identifying true positives across all classes.
@@
Figure 5: Macro recall across all 5 models. OpenVLA performs better when considering macro recall compared to macro precision, indicating a high number of false positives. GPT 4o shows lower macro recall than precision indicating biased performance towards specific minority classes. GPT 4.1 and Pi0 Base show relatively less biased and moderate performance, whereas Pi0 FAST showed consistent low recall.
@@
The analysis of invalid predictions highlighted significant disparities. OpenVLA, by design, produced zero invalid predictions, ensuring actions always fell within the valid action space. As seen in Figure 6 GPT 4o and Pi0 FAST frequently generated invalid outputs, with some datasets reaching invalid prediction rates above 80 percent. Conversely, GPT 4.1 and Pi0 Base demonstrated notably lower invalid prediction rates, underscoring their relatively stronger capability to adhere to the valid action space constraints.
@@
Figure 6: Percentage invalids across all 5 models. Invalids refer to model predictions that are not valid actions in the subdataset’s action space. Pi0 FAST and GPT 4o struggle to produce valid actions irrespective of the subdataset
@@ IV-A2 Outlier Influence and Majority Prediction Performance
As seen in Figures 7 and 8 the normalized Brier MAE and quantile-filtered analysis indicated consistent error trends when compared to regular Brier MAE scores across models and datasets, demonstrating minimal reliance on outlier predictions. Given the already high median errors, the influence of outliers was negligible.
Figure 7: Normalized Brier MAE across 4 models - GPT 4o, OpenVLA, GPT 4.1, Pi0 FAST. Follows trends of regular Brier MAE indicating minimal reliance on outliers
Figure 8: Normalized Quantile Filtered Brier MAE across 4 models. Follows trends of regular and Normalized Brier MAE scores indicating minimal reliance on extreme predictions
Further investigation using the max relative Brier MAE metric showed minimal variance in model performance, the scores staying consistently close to 1.0. This finding, seen in Figure 9, suggests limited deviation between median and extreme error values, primarily due to the universally high error levels across all models, which reduce the relative significance of outlier predictions.
@@ IV-B1 Dataset-Specific Performance Across Models
OpenVLA consistently achieved the highest macro-recall among the evaluated models across nearly all 16 Procgen subdatasets, typically ranging from 9 percent to 12 percent. It performed notably better on simpler tasks, achieving a maximum recall of approximately 12.5 percent on the Maze dataset, characterized by straightforward navigation and collection actions without the complexity of additional special actions. Conversely, OpenVLA’s recall dropped to its lowest ( 8 percent) on the Ninja dataset, which involves multiple directional special ”fire” actions, indicating difficulty handling tasks requiring timely execution of specialized actions.
@@
GPT-4.1 and Pi0 Base occupied a middle performance tier, displaying macro-recall values roughly between 6 percent and 11 percent. GPT-4.1 generally outperformed GPT-4o across most datasets and showed superior performance compared to Pi0 Base on simpler datasets like Leaper, Climber, Jumper, Maze and Heist. Conversely, GPT-4o exhibited significantly weaker performance than Pi0 Base on datasets involving timely and sparse execution of special actions such as Plunder, Fruitbot, and Starpilot. These three datasets involve ”fire” actions, indicating GPT-4o’s difficulty in handling timely execution tasks. Specifically, GPT-4.1 achieved macro-recall scores around 11 percent on simpler datasets including Leaper, Climber, Jumper, Bigfish, Chaser, Coinrun, Miner, Maze, and Heist. In contrast, its performance diminished to around 7 percent on more challenging environments such as Plunder, Starpilot, Bossfight, and Ninja. GPT-4o showed the highest recalls around 8.5 percent on datasets with simpler movement-only actions (Leaper, Climber, Jumper, Coinrun, Maze, Heist), but performed poorly (around 3 percent) on more complex tasks like Chaser and Bossfight, and extremely poorly (below 2.5 percent) on Fruitbot and Starpilot, both featuring special actions.
@@
Pi0 Base achieved macro recalls typically between 6 percent and 10 percent. Its best performance (around 10 percent) appeared on datasets such as Leaper, Plunder, Jumper, Bigfish, Chaser, Coinrun, and Miner. Conversely, macro recall dropped to approximately 6 percent on datasets involving special actions like Dodgeball, Fruitbot, Bossfight, and Ninja.
@@
Pi0 FAST exhibited the lowest macro recall among the evaluated models, frequently below 6 percent, and occasionally as low as 2 percent on datasets including Leaper, Bigfish, Coinrun, Ninja, Miner, Maze, and Heist. However, Pi0 FAST displayed anomalously higher recalls ( 10 percent) on datasets such as Starpilot and Bossfight. Despite these datasets involving special actions, the actions are executed frequently rather than sparsely, aligning better with Pi0 FAST’s predictions and ground truth distributions.
@@
In general, our findings highlight a clear difficulty across models in accurately handling datasets that require complex, sparse, or timely executed actions, particularly when special non-movement actions are involved. The macro recall performance appears directly correlated to how closely model predictions distribution aligns with the ground truth action distribution.
@@ IV-B2 Action-Class Specific Performance Across Models
A detailed analysis of recall performance by individual action classes offers valuable insights into model-specific strengths, weaknesses, and intrinsic biases in predicting discrete actions. As seen in Figure 10 action classes 4 through 10 generally displayed higher recall scores and greater variance among models, indicating varying degrees of predictive confidence. In contrast, action classes 0 to 3 and 11 to 12 consistently yielded low recall values across all models, demonstrating limited predictive capability.
@@
GPT-4o exhibited notable recall performance on action class 7 (”Right” movement), achieving a high recall of approximately 0.22. This suggests a directional bias likely influenced by the prevalent rightward progression in many Procgen tasks. Conversely, GPT-4o struggled significantly with actions 4, 10, 11, and 12, displaying near-zero recall, highlighting clear limitations in handling certain peripheral or specialized actions.
@@
OpenVLA predominantly favored action class 4 (”Do Nothing”), recording the highest recall ( 0.32). This bias arises from a combined influence of its decoding mechanism and Procgen action space sizes. Most Procgen datasets have an action space of size 9 or 10, and due to unnormalization of OpenVLA’s predicted actions using the dataset statistics, 4 can correspond to 0 in the normalized range which is the center of OpenVLA’s output distribution. However, OpenVLA encountered considerable difficulty predicting peripheral classes (0, 1, 8, 10, and 11), with recall scores as low as 0.03, indicating a substantial predictive bias toward central actions.
@@
Pi0 Base similarly gravitated toward central action classes (4, 5, 6, and 7) due to its normalization-unormalization decoding approach, resulting in moderate recall values (  0.12 - 0.16). Its recall sharply declined for peripheral and special action classes (0, 1, 2, and 12), dropping to approximately 0.02. Compared to OpenVLA, Pi0 Base’s diffusion-based approach led to less concentrated predictions, distributing recall more evenly but still inadequately across peripheral classes.
@@
Pi0 FAST displayed a pronounced bias towards higher-numbered action classes (8, 9, and 10), with recall scores around 0.10. This bias reflects its difficulty generalizing to OOD data, often producing higher normalized action predictions. Conversely, Pi0 FAST severely underperformed on lower-numbered and mid-range actions (0, 2, 3, 7, and 11), frequently achieving near-zero recall, emphasizing its limitations in balanced action class predictions.
@@
GPT-4.1 demonstrated a distinct bias toward action class 7 (”Right” movement), achieving exceptionally high recall ( 0.40), consistent with the directional tendencies of many Procgen environments. It also showed moderate recall ( 0.20) for special action class 9. However, GPT-4.1 struggled notably with several other action classes, particularly actions 0, 1, 4, 10, and 11, consistently resulting in minimal recall ( 0.01).
@@ IV-B3 Models’ Prediction Collapse to Few Action Classes
A consistent observation across all evaluated models was their tendency to collapse predictions predominantly into a limited subset of action classes. Analysis of union confusion matrices as seen in Figure 11 aggregated over all 16 Procgen subdatasets highlighted clear default prediction behaviors.
@@
Lists are easy to create:
  • OpenVLA frequently defaulted to predicting action classes 4 (”Do Nothing”) and 6 (”RIGHT + DOWN”).
  • Pi0 Base predominantly predicted action classes clustered between 4 to 7 (”Do Nothing,” ”UP,” ”RIGHT + DOWN,” ”RIGHT”) and occasionally between 5 to 8 (”UP,” ”RIGHT + DOWN,” ”RIGHT,” ”RIGHT + UP”).
  • Pi0 FAST consistently favored action classes 6 (”RIGHT + DOWN”) and 8 (”RIGHT + UP”).
  • GPT-4o exhibited a strong bias towards action class 7 (”RIGHT”).
  • GPT-4.1 frequently defaulted to action classes 7 (”RIGHT”) and 2 (”LEFT + UP”).
Figure 11: Union confusion matrices depicting frequency of class predictions and ground truth matches. Order from top to bottom - GPT 4o, OpenVLA, Pi0 Base, Pi0 FAST, GPT 4.1
Notably, as seen in Figure 12, across all models and subdatasets, action class 4 (”Do Nothing”) exhibited significantly higher precision compared to other classes. This elevated precision for class 4 likely originates from its role as a default assignment for expert actions that fall outside the environment-specific action spaces when creating ground truth labels, thus inflating its representation in datasets.
Figure 12: Class-wise precision averaged across all subdatasets for all 5 models. Precision is significantly higher for 4 across all models
For autoregressive models like GPT-4o and GPT-4.1, the frequent default to action class 4 may also reflect insufficient understanding of the prompt and task, causing uncertainty in predictions and defaulting to ”Do Nothing.” Additionally, the observed strong preference for ”RIGHT” (action class 7) aligns with the prevalent directional progression in many Procgen environments.
In contrast, OpenVLA, Pi0 Base, and Pi0 FAST primarily defaulted towards actions situated near the center of their normalized prediction range (typically zero). Upon unnormalization, these central predictions commonly mapped to action class 4 due to rounding effects and the typical size (<<<= 10 actions) of Procgen datasets.

### v2

@@ IV-A1 Comparing General Performance
In evaluating the generalization capability of GPT 4o, GPT 4.1, OpenVLA, Pi0 Base, and Pi0 FAST, we observed universally poor performance on zero-shot out-of-distribution (OOD) tasks. The analysis using various performance metrics underscored significant limitations across these models.
As seen in Figure 2, the Brier Mean Absolute Error (MAE) scores revealed poor probability calibration, as the scores approached the maximum value of 2 across all models and datasets. Specifically, GPT 4o, GPT 4.1, and Pi0 FAST consistently scored above 1.70, while OpenVLA generally exceeded 1.50. The subdatasets on which models exhibited the weakest performance were Starpilot for GPT 4o, Leaper for GPT 4.1, Miner for Pi0 FAST, and Ninja for OpenVLA.
Figure 2: Brier Mean Absolute Error scores across 4 models - GPT 4o, OpenVLA, GPT 4.1, and Pi0 FAST. Pi0 Base is a diffusion-based model and can not be evaluated using Brier MAE due to a lack of logits in its inference architecture. All models display Brier MAE close to 2, indicating poor performance.
Analyzing micro-level metrics, precision values were low across all models and subdatasets. OpenVLA achieved the highest micro precision at approximately 27% on the Coinrun dataset, whereas GPT 4o showed the lowest precision at 1% on Starpilot. Pi0 FAST reached a minimal precision of around 1% on Heist. Overall, as can be inferred from Figure 3, the performance hierarchy based on micro precision was: OpenVLA $>$ Pi0 Base $>$ GPT 4.1 $>$ GPT 4o $>$ Pi0 FAST.
@@
When considering micro precision without invalid predictions, as seen in Figure 4, significant improvements were noted for GPT 4o and Pi0 FAST, highlighting their potential if constrained or further fine-tuned to produce valid action outputs. GPT 4o’s precision notably improved, surpassing Pi0 Base in a few subdatasets such as Coinrun, Bossfight, and Dodgeball. Nevertheless, the adjusted performance order still positioned OpenVLA at the top, followed by Pi0 Base, Pi0 FAST, GPT 4o, and GPT 4.1.
Figure 4: Micro precision across all 5 models - GPT 4o, OpenVLA, Pi0 Base, Pi0 FAST, and GPT 4.1 without invalids. When considering all 16 subdatasets, GPT 4.1 overall showcases the worst performance, and OpenVLA the best performance.
@@
Macro precision analysis revealed that GPT 4o, OpenVLA, and Pi0 Base maintained similar performance levels across most datasets. As seen in Figure 5, conversely, GPT 4.1 and Pi0 FAST demonstrated significant variability, implying distinct preferences for specific task types. Despite these observations, macro precision remained generally low across all models, with a maximum slightly above 15%. Importantly, comparing its macro and micro precision performances, it is clear that OpenVLA’s performance heavily favored the majority classes, while GPT 4o exhibited relatively stronger performance on minority classes.
@@
Figure 5: Macro precision across all 5 models. GPT 4.1 and Pi0 FAST show variability and preference towards specific subdatasets. OpenVLA shows reliability on the majority classes, whereas GPT 4o displays relatively stronger performance on minority classes.
@@
Macro recall, as seen in Figure 6, highlighted OpenVLA’s aggressive prediction strategy, resulting in consistently high recall but lower precision due to numerous false positives. GPT 4o, despite higher macro precision, showed substantially lower recall, indicating pronounced class biases towards a few classes, and a potentially overly conservative nature about predicting certain classes compared to other classes. GPT 4.1 exhibited a relatively balanced approach, achieving both recall and precision, thus indicating less bias towards specific classes. Pi0 Base displayed a moderately balanced performance without leading in any subdataset, whereas Pi0 FAST consistently showed low recall, reflecting challenges in identifying true positives across all classes.
@@
Figure 6: Macro recall across all 5 models. OpenVLA performs better when considering macro recall compared to macro precision, indicating a high number of false positives. GPT 4o shows lower macro recall than precision, indicating biased performance towards specific minority classes. GPT 4.1 and Pi0 Base show relatively less biased and moderate performance, whereas Pi0 FAST showed consistent low recall.
@@
The analysis of invalid predictions highlighted significant disparities. OpenVLA, by design, produced zero invalid predictions, ensuring actions always fell within the valid action space. As seen in Figure 7, GPT 4o and Pi0 FAST frequently generated invalid outputs, with some datasets reaching invalid prediction rates above 80%.
@@
Figure 7: Percentage invalids across all 5 models. Invalids refer to model predictions that are not valid actions in the subdataset’s action space. Pi0 FAST and GPT 4o struggle to produce valid actions irrespective of the subdataset.
@@ IV-A2 Outlier Influence and Majority Prediction Performance
As seen in Figures 8 and 9, the normalized Brier MAE and quantile-filtered analysis indicated consistent error trends when compared to regular Brier MAE scores across models and datasets, demonstrating minimal reliance on outlier predictions. Given the already high median errors, the influence of outliers was negligible.
Figure 8: Normalized Brier MAE across 4 models - GPT 4o, OpenVLA, GPT 4.1, Pi0 FAST. It follows the trends of regular Brier MAE, indicating minimal reliance on outliers.
Figure 9: Normalized Quantile Filtered Brier MAE across 4 models. It follows trends of regular and Normalized Brier MAE scores, indicating minimal reliance on extreme predictions.
Further investigation using the max relative Brier MAE metric showed minimal variance in model performance, with the scores staying consistently close to 1.0. This finding, seen in Figure 10, suggests limited deviation between median and extreme error values, primarily due to the universally high error levels across all models reducing the relative significance of outlier predictions.
@@ IV-B1 Dataset-Specific Performance Across Models
Figure 11: Distribution of model macro recall across subdatasets. The black and red lines indicate the mean and median, respectively.
@@
OpenVLA consistently achieved the highest macro-recall among the evaluated models across nearly all 16 Procgen subdatasets, typically ranging from 9% to 12% as shown in 6 and 11. It performed notably better on simpler tasks, achieving a maximum recall of approximately 12.5% on the Maze dataset, characterized by straightforward navigation and collection actions without the complexity of additional special actions. Conversely, OpenVLA’s recall dropped to its lowest ( 8 percent) on the Ninja dataset, which involves multiple directional special “fire” actions, indicating difficulty handling tasks requiring timely execution of specialized actions. This could also indicate increased unfamiliarity with the action type since movement actions may be more similar to the model’s training environment than special actions.
@@
GPT-4 Series occupied a middle performance tier, displaying macro-recall values roughly between 6% and 11% as seen in 1 and 11. GPT-4.1 generally outperformed GPT-4o across most datasets. Notably, GPT-4.1 showed superior performance when compared to Pi0 Base on simpler datasets like Leaper, Climber, Jumper, Maze, and Heist. Specifically, GPT-4.1 achieved macro-recall scores of around 11% on simpler datasets like Leaper, Climber, Jumper, Bigfish, Chaser, Coinrun, Miner, Maze, and Heist. Its performance diminished to around 7% on more challenging environments such as Plunder, Starpilot, Bossfight, and Ninja.
Conversely, GPT-4o exhibited significantly weaker performance than Pi0 Base on datasets involving timely and sparse execution of special actions. This was particularly evident in datasets involving “fire” actions, such as Plunder, Fruitbot, and Starpilot, indicating GPT-4o’s difficulty managing tasks requiring precise timing. While GPT-4o showed the highest recalls around 8.5% on datasets with simpler movement-only actions (Leaper, Climber, Jumper, Coinrun, Maze, Heist), it performed significantly worse on more complex tasks. Recall dropped to around 3% on more complex tasks like Chaser and Bossfight, and below 2.5% on Fruitbot and Starpilot, both of which feature special actions.
@@
Pi0 Base achieved macro recalls typically between 6% and 10%, as seen in 1 and 11. Its best performance (around 10 percent) appeared on relatively simpler datasets such as Leaper, Plunder, Jumper, Bigfish, Chaser, Coinrun, and Miner. Conversely, macro recall dropped to approximately 6% on datasets involving special actions like Dodgeball, Fruitbot, Bossfight, and Ninja.
@@
Pi0 FAST exhibited the lowest macro recall among the evaluated models, frequently below 6 percent, and occasionally as low as 2% on datasets including Leaper, Bigfish, Coinrun, Ninja, Miner, Maze, and Heist as seen in 1 and 11. However, Pi0 FAST displayed anomalously higher recalls ( 10 percent) on datasets such as Starpilot and Bossfight. Despite these datasets involving special actions, the actions are executed frequently rather than sparsely, aligning better with Pi0 FAST’s predictions and ground truth distributions.
@@
In general, our findings highlight a clear difficulty across models in accurately handling datasets that require complex, sparse, or timely executed actions, particularly when special non-movement actions are involved. The macro recall performance appears directly correlated to how closely the distribution of model predictions aligns with the distribution of the ground truth actions.
@@ IV-B2 Action-Class Specific Performance Across Models
A detailed analysis of recall performance by individual action classes offers valuable insights into model-specific strengths, weaknesses, and intrinsic biases in predicting discrete actions. As seen in Figure 12, action classes 4 through 10 generally displayed higher recall scores and greater variance among models, indicating varying degrees of predictive confidence. In contrast, action classes 0 to 3 and 11 to 12 consistently yielded low recall values across all models, demonstrating limited predictive capability.
Figure 13: Model action class recall vs. action class frequency
When comparing the model action class recall against action class frequency in Figure 13, OpenVLA exhibits the strongest positive correlation, attaining the highest recall on the most prevalent classes, while Pi0 base shows little correlation. Both GPT-4o and GPT-4.1 display slightly negative trends, indicating more struggle with common classes than with rarer ones. Pi0 Fast’s recall remains largely constant across action class frequencies, indicating uniform yet low performance.
@@
GPT-4o exhibited notable recall performance on action class 7(“Right” movement), achieving a high recall of approximately 0.22 as seen in 12. This suggests a directional bias likely influenced by the prevalent rightward progression in many Procgen tasks. Conversely, GPT-4o struggled significantly with actions 4, 10, 11, and 12, displaying near-zero recall, highlighting clear limitations in handling certain peripheral or specialized actions. Specifically, the difference between the most frequent action classes in predicted actions and ground truth actions can be seen in Figure 14.
@@
OpenVLA predominantly favored action class 4 (“Do Nothing”) as seen in 15, recording the highest recall ( 0.32) as seen in 12. This bias arises from a combined influence of its decoding mechanism and Procgen action space sizes. Most Procgen datasets have an action space of size 9 or 10, and due to unnormalization of OpenVLA’s predicted actions using the dataset statistics, 4 can correspond to 0 in the normalized range, which is the center of OpenVLA’s output distribution. OpenVLA displays considerable difficulty predicting peripheral classes (0, 1, 8, 10, and 11), with recall scores as low as 0.03, indicating a substantial predictive bias toward central actions.
@@
Pi0 Base similarly gravitated toward central action classes (4, 5, 6, and 7) as seen in 16 due to its normalization-unormalization decoding approach, resulting in moderate recall values (  0.12 - 0.16) as seen in 12. Its recall sharply declined for peripheral and special action classes (0, 1, 2, and 12), dropping to approximately 0.02. Compared to OpenVLA, Pi0 Base’s diffusion-based approach led to less concentrated predictions, distributing recall more evenly but still inadequately across peripheral classes.
@@
Pi0 FAST displayed a pronounced bias towards action classes (1, 5, 8, and 9) as seen in 17, with action percentage between 10% and 26%. This bias reflects its difficulty generalizing to OOD data, often producing higher normalized action predictions. Conversely, Pi0 FAST severely underperformed on actions between the above prevalent classes, such as 0, 2, 3, 7, and 11, frequently achieving near-zero recall as seen in 12, emphasizing its limitations in balanced action class predictions.
@@
GPT-4.1 demonstrated a distinct bias toward action class 7 (“Right” movement) as seen in 18, achieving exceptionally high recall ( 0.40) as seen in 12, consistent with the directional tendencies of many Procgen environments. It also showed moderate recall ( 0.20) for special action class 9. However, GPT-4.1 struggled notably with several other action classes, particularly actions 0, 1, 4, 10, and 11, consistently resulting in minimal recall ( 0.01).
@@
Figure 14: GPT4o prediction vs. ground truth distributions.
Figure 15: OpenVLA prediction vs. ground truth distributions
Figure 16: Pi0 base prediction vs. ground truth distributions
Figure 17: Pi0 fast prediction vs. ground truth distributions
Figure 18: GPT-4.1 prediction vs. ground truth distributions
@@ IV-B3 Models’ Prediction Collapse to Few Action Classes
A consistent observation across all evaluated models was their tendency to collapse predictions predominantly into a limited subset of action classes. Analysis of union confusion matrices as seen in Figure 19 and action distributions 14, 15, 16, 17, and 18 aggregated over all 16 Procgen subdatasets highlighted clear default prediction behaviors.
@@
  • OpenVLA frequently defaulted to predicting action classes 4 (“Do Nothing”), 5 (“UP”), and 6 (“RIGHT + DOWN”).
  • Pi0 Base predominantly predicted action classes clustered between 4 to 7 (“Do Nothing”, “UP”, “RIGHT + DOWN”, “RIGHT”) and the range shifts to between 5 to 8 (“UP”, “RIGHT + DOWN”, “RIGHT”, “RIGHT + UP”) when the ground truth actions are between 9 and 12.
  • Pi0 FAST consistently favored action classes 6 (“RIGHT + DOWN”) and 8 (“RIGHT + UP”).
  • GPT-4o exhibited a strong bias towards action class 7 (“RIGHT”).
  • GPT-4.1 frequently defaulted to action classes 7 (“RIGHT”) and 2 (“LEFT + UP”).
Figure 19: Union confusion matrices depicting frequency of class predictions and ground truth matches. Order from top to bottom - GPT 4o, OpenVLA, Pi0 Base, Pi0 FAST, GPT 4.1.
Notably, as seen in Figure 20, across all models and subdatasets, action class 4 (“Do Nothing”) exhibited significantly higher precision compared to other classes. We speculate that this elevated precision for class 4 originates from its role as a default assignment for expert actions that fell outside the environment-specific action spaces during the generation of ground truth labels, thus inflating its representation.
Figure 20: Class-wise precision averaged across all subdatasets for all 5 models. Precision is significantly higher for 4 across all models.
For VLMs like GPT-4o and GPT-4.1, the frequent choice of action class 4 may also reflect insufficient understanding of the task, causing it to be uncertain in its predictions and default to “Do Nothing”. Additionally, the observed strong preference for “RIGHT” (action class 7) aligns with the prevalent directional progression in many Procgen environments.
In contrast, OpenVLA, Pi0 Base, and Pi0 FAST primarily defaulted towards actions situated near the center of their normalized prediction range (typically zero). These central predictions commonly map to action class 4 due to rounding effects and the typical size (<<<= 10 actions) of Procgen datasets.